# A Novel Benchmark for Few-Shot Semantic Segmentation in the Era of Foundation Models

### v1

@@ 4.1 Training procedure
4.2 Main results
In this section, we aim to identify the optimal combination of pretrained models and adaptation methods based on mIoU. Table 1 presents the average performance on the three considered datasets for all possible combinations of pretrained models and adaptation methods, along with their standard deviation, in the context of 1-shot problems.
From this initial experiment, several conclusions can be drawn:
  • SAM encoder exhibits strong performance on average. It outperforms CLIP, MAE, and ResNet on both Cityscapes and PPD. However, it yields comparatively poor results on COCO (refer to Appendix Table 2). This difference is likely attributed to our decision to retain the SAM image encoder while discarding the original decoder, which was jointly trained with the image and prompt encoders.
  • The multilayer method outperforms the linear approach for most models, indicating that including earlier feature maps, which are more locally focused [25, 14], provides the decoder with greater segmentation flexibility. Furthermore, it exhibits a lower standard deviation, indicating greater stability across different runs.
  • Despite variations in the number of training parameters (less than 1% for LoRA, around 2% for SVF, and 100% for finetune), the finetuning methods (SVF, LoRA, and Finetune) yield comparable results.
@@ 5.2 Semantic Segmentation and Few-Shot Semantic Segmentation
Few-shot semantic segmentation is not a recent topic. Earlier works, such as [2] and [39], explored multiple benchmarks, including PASCAL-$5^{i}$ and COCO-$20^{i}$, reminiscent of in-domain benchmarks used in few-shot classification. Subsequent research ventured into using pretrained models on ImageNet [40], resembling the Meta-dataset approach in few-shot classification. Notably, most of these studies assumed binary labelling for the few-shots, emphasizing a single object of interest.
In contrast, our approach considers that the provided few-shots are fully labelled, and we leverage pretrained foundation models, not limited to those trained solely on ImageNet. This approach aligns more closely with contemporary industrial requirements, focusing on the utilization of the most efficient available models.
@@ 6 Conclusion
In conclusion, DINO V2 emerges as the top-performing backbone model, consistently outperforming other foundation models. While SAM exhibits strong overall results, it presents unexpected limitations on certain datasets. Surprisingly, our experiments underscore the competitive advantages of straightforward methods such as simple linear heads trained on the few given shots, challenging the necessity of complex procedures.
This work not only presents a comparison of techniques but also a practical benchmark. We hope this benchmark will contribute to improving few-shot semantic segmentation methods and will provide a solid ground for comparison purposes.

### v2

@@ 5.2 What Works The Best?
We then vary the number of shots to 2, 5 and 10. Table 2 shows the consistency of the results with 1-shot, where:
  • DINO V2 outperforms other models by a big margin.
  • SAM still yields strong performance on average, but performs poorly on COCO (refer to appendix Tables 101112).
  • The multilayer method outperforms the linear approach for most models.
  • Fine-tuning methods yield comparable results.
In addition, we can see that:
  • DINO V2 scales better than other models with more shots.
  • Fine-tuning scales better than LoRA and SVF and gives better results on average on the 10-shot setting, this is due to the fact that we suffer less from over-fitting when having more shots.
Table 2: mIoU averaged over our three considered datasets in $k$-shot semantic segmentation for $k\in\{2,5,10\}$. Bold numbers correspond to the best performance across extractors and underline numbers to the best performance across adaptation methods.
Linear Multilayer SVF LoRA Fine-tuning
[…]
DINO V2 61.20$\pm$00.29 59.21$\pm$00.83 64.10$\pm$00.83 63.73$\pm$01.28 64.35$\pm$00.33
CLIP 39.08$\pm$03.45 43.46$\pm$00.70 45.59$\pm$01.94 42.72$\pm$02.97 35.68$\pm$07.11
MAE 36.90$\pm$04.15 38.73$\pm$02.39 40.36$\pm$03.83 40.37$\pm$04.27 39.15$\pm$03.50
FCN-ResNet50 40.98$\pm$01.96 47.36$\pm$00.85 48.66$\pm$00.92 46.55$\pm$02.36 48.81$\pm$00.32
[…]
DINO V2 66.47$\pm$00.60 65.89$\pm$00.63 68.64$\pm$00.88 67.54$\pm$00.20 68.82$\pm$01.40
CLIP 46.30$\pm$00.29 48.67$\pm$00.61 51.56$\pm$00.22 49.00$\pm$00.39 43.60$\pm$00.76
MAE 44.36$\pm$00.23 45.02$\pm$00.15 46.20$\pm$00.92 47.02$\pm$00.16 46.68$\pm$00.34
FCN-ResNet50 47.02$\pm$00.48 49.70$\pm$00.27 51.58$\pm$00.03 51.40$\pm$00.28 51.51$\pm$00.26
[…]
DINO V2 69.86$\pm$00.23 69.95$\pm$00.12 69.97$\pm$00.69 69.25$\pm$00.25 71.07$\pm$01.35
CLIP 50.14$\pm$00.41 52.49$\pm$00.41 53.64$\pm$00.77 51.92$\pm$00.41 47.62$\pm$03.36
MAE 47.56$\pm$00.41 47.97$\pm$00.28 46.73$\pm$00.17 42.94$\pm$03.07 48.63$\pm$00.30
FCN-ResNet50 48.69$\pm$00.69 51.98$\pm$00.25 53.04$\pm$00.31 53.02$\pm$00.25 52.86$\pm$00.55
5.3 Practical Considerations
Previous values were obtained by selecting the best learning rates on the validation sets. Instead, here we study the effect of transfering the learning rate from a dataset to another one on the mIoU. The learning rate values we explored are $10^{-2}$, $10^{-3}$, $10^{-4}$, $10^{-5}$ and $10^{-6}$. Table 3 shows the drop of mIoU when we train a pair (model, method) on dataset A with the best learning rate for dataset B. We can see that the learning rate is relatively consistent between the 3 datasets and especially between Cityscapes and COCO for a given model and a given method. even though the optimal values are not exactly the same between these datasets, the variation of performance is low and applying the optimal learning of one dataset on the other does not cause a big drop of performance, so we argue that these values could serve as a good baseline for a new unseen dataset. The higher drops occur when we either take the optimal learning rate for PPD or when we test an optimal learning rate of Cityscapes or COCO on PPD. This is mainly due to the fact that the support set of PPD is significantly smaller than that of the other considered datasets.
Table 3: Drop in mIoU when using the best learning rate corresponding to the dataset on the line (source) when training on the dataset on the column (target), instead of using the target validation set to find the best learning rate.
SVF LoRA Fine-tuning
Cityscapes COCO PPD Cityscapes COCO PPD Cityscapes COCO PPD
[…]
COCO 0 - 0 0 - 0 0.94 - 3.01
PPD 0 0 - 0 0 - 3.38 3.83 -
[…]
COCO 0 - 0 0 - 2.00 0 - 26.3
PPD 0 0 - 0.88 0.88 - 1.17 1.17 -
[…]
COCO 0 - 1.09 0.24 - 0 0 - 0
PPD 1.91 1.91 - 1.53 0 - 0 0 -
[…]
COCO 0.28 - 0 0.87 - 0 0.65 - 0.65
PPD 1.41 0 - 1.47 0 - 0 1.53 -
[…]
COCO 0.88 - 4.07 0.66 - 0.66 1.55 - 1.55
PPD 8.46 18.01 - 0 7.85 - 0 12.13 -
5.4 Individual Factors
In the next series of experiments, we study the effect of various elements that could influence the performance of the models. Namely, these include the model size, the underlying architecture, the pretrained dataset and the pretraining method. By studying these individual components, we seek to provide a more granular understanding of what drives model performance.
5.4.1 Input Resolution.
We explore how input resolution impacts model performance, a critical factor given our training setup. Specifically, CLIP and MAE were trained at a resolution of 224x224, while other models were trained with a higher resolution of 1024x1024. This discrepancy is due to the fixed input size requirements of CLIP and MAE’s positional embeddings. To address this, we adapted the positional embeddings matrix for variable resolutions, a common practice for ViT models. Our findings, illustrated in Figure 3, reveal a divergence in performance trends: increasing resolution degrades results for CLIP and MAE but improves them for DINO V2. This can be attributed to DINO V2’s training across multiple resolutions, unlike CLIP and MAE. Consequently, this suggests using 1024x1024 input images on DINO V2 whereas 224x224 input images on CLIP and MAE is fair for performance comparison.
224512768102420304050607080Input ResolutionmIoUDINO V2CLIPMAE
Figure 3: Impact of the input resolution on the obtained mIoU.
5.4.2 Model Size.
Table 4 demonstrates that a larger model size does not translate to a big boost of performance. Specifically, for SAM and DINO V2 with the linear method, the base model (ViT-B) yields, on average, competitive performance, despite having 3 times less parameters than the (ViT-L) models. In the case of ResNet, moving from ResNet-50 to ResNet-101 shows a performance boost, though the increase is modest. For the LoRA method, fine-tuning of a subset of parameters of the encoder yields better performance with bigger models. These observations partially accounts for the performance discrepancy between SAM, DINO V2, and ResNet50, highlighting the influence of model size differences.
Table 4: Effect of model size on mIoU.
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
SAM (ViT-B) 35.72$\pm$01.50 03.13$\pm$00.09 83.14$\pm$10.36 38.50$\pm$00.12 05.85$\pm$00.13 84.94$\pm$10.11
SAM (ViT-L) 35.00$\pm$01.29 03.43$\pm$00.07 72.95$\pm$09.05 40.70$\pm$00.45 07.31$\pm$00.37 79.14$\pm$10.75
SAM (ViT-H) 34.64$\pm$01.36 03.38$\pm$00.11 83.34$\pm$04.02 40.78$\pm$00.57 07.46$\pm$00.29 85.73$\pm$08.42
DINO V2 (Vit-S) 41.16$\pm$02.39 20.62$\pm$00.38 87.99$\pm$05.58 44.75$\pm$00.71 22.72$\pm$00.71 86.55$\pm$06.08
DINO V2 (Vit-B) 48.80$\pm$01.82 23.24$\pm$00.38 92.31$\pm$01.84 54.35$\pm$01.55 28.99$\pm$01.33 89.67$\pm$06.43
DINO V2 (Vit-L) 49.36$\pm$01.61 23.50$\pm$00.33 87.87$\pm$07.30 54.79$\pm$00.64 31.54$\pm$00.81 87.46$\pm$07.97
FCN-ResNet50 32.39$\pm$01.79 10.08$\pm$00.32 70.86$\pm$05.91 35.74$\pm$01.41 11.71$\pm$00.24 76.91$\pm$10.29
FCN-ResNet101 32.97$\pm$01.76 10.94$\pm$00.16 69.70$\pm$12.98 29.38$\pm$07.03 12.23$\pm$00.57 81.79$\pm$11.40
5.4.3 Training Dataset.
To examine the impact of the training dataset on model performance, we compared DINO V1 and DINO V2, noting that the primary distinction between them lies in their training datasets. DINO V1 was trained on ImageNet-1K, whereas DINO V2 utilized a much larger dataset comprising 142M curated and uncurated images. It is important to mention that another difference between the two models is the employed loss function, although the focus of this analysis is primarily on the effect of dataset size and composition. Table 5 illustrates that DINO V2 significantly surpasses DINO V1 in performance, with DINO V1 exhibiting notably weak results on the COCO dataset. Showing that having a bigger training dataset is beneficial for the downstream task performance on few-shot semantic segmentation.
Table 5: Effect of the training dataset on mIoU
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
DINO V1 29.03$\pm$01.08 05.92$\pm$00.39 82.14$\pm$10.04 32.53$\pm$00.97 06.89$\pm$00.44 82.99$\pm$09.35
DINO V2 48.80$\pm$01.82 23.24$\pm$00.38 92.31$\pm$01.84 54.35$\pm$01.55 28.99$\pm$01.33 89.67$\pm$06.43
5.4.4 Architecture.
To assess the influence of model architecture, we analyzed the performance differences between DINO V1 (ViT-S) and its ResNet-50 variant, alongside comparing the CLIP (ViT-B) version against the CLIP ResNet-50 version. It is noteworthy that for CLIP, the ViT-B model is larger than its ResNet counterpart, an aspect previously discussed in terms of model size impact. According to Table 6, ViT architectures demonstrate a clear performance advantage over their ResNet equivalents. This trend underscores the potential superiority of ViT models in adapting to various tasks.
Table 6: Effect of model architecture on mIoU
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
DINO V1 (ViT-S) 26.61$\pm$00.53 04.86$\pm$00.23 81.86$\pm$08.19 29.90$\pm$00.95 05.82$\pm$00.34 79.70$\pm$09.03
DINO V1 (ResNet-50) 23.51$\pm$01.40 03.42$\pm$00.05 58.98$\pm$00.61 26.98$\pm$01.28 03.78$\pm$00.04 77.35$\pm$10.50
CLIP (ViT-B) 23.83$\pm$01.51 11.32$\pm$00.31 72.49$\pm$10.78 27.95$\pm$01.62 13.62$\pm$00.78 73.13$\pm$11.70
CLIP (ResNet-101) 20.22$\pm$01.57 02.61$\pm$00.06 69.00$\pm$06.03 08.96$\pm$00.11 00.91$\pm$00.03 67.40$\pm$11.91
5.4.5 Training Method.
To investigate the effect of different self-supervised learning (SSL) methods on model performance, we conducted a comparison between DINO V1 and MAE. Both models were trained on the same dataset (ImageNet-1K), and share identical architecture and size (ViT-B with 86M parameters), making the SSL method the primary variable. The comparison in Table 7 reveals that DINO V1 surpasses MAE on the Cityscapes and PPD datasets, whereas MAE demonstrates superior performance on COCO. This mixed outcome indicates that there is no definitive answer as to which SSL training approach offers the best adaptability for few-shot segmentation tasks.
Table 7: Effect of the training method on mIoU
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
DINO V1 29.03$\pm$01.08 05.92$\pm$00.39 82.14$\pm$10.04 32.53$\pm$00.97 06.89$\pm$00.44 82.99$\pm$09.35
MAE 23.63$\pm$00.89 08.54$\pm$00.46 74.34$\pm$10.88 26.10$\pm$00.26 10.13$\pm$00.42 72.87$\pm$12.83
5.4.6 Registers.
In the study by [5], it is demonstrated that DINO V2, unlike its predecessor DINO V1, presents artifacts within its feature maps. This phenomenon is attributed to the presence of tokens with unusually high norms, which are integral to the model’s internal computations. To address this issue, the authors introduced a modified version of DINO V2, incorporating additional tokens termed “Registers” during the training phase, which are subsequently omitted during inference. As evidenced in Table 8, this innovative approach significantly enhances DINO V2’s performance on the COCO dataset, showcasing the potential benefits of Registers during training for few-shot segmentation tasks.
Table 8: Effect of training with registers on mIoU
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
DINO V2 w/o Registers 48.80$\pm$01.82 23.24$\pm$00.38 92.31$\pm$01.84 54.35$\pm$01.55 28.99$\pm$01.33 89.67$\pm$06.43
DINO V2 w/ Registers 48.06$\pm$02.46 27.42$\pm$00.72 91.58$\pm$02.41 54.08$\pm$00.44 34.02$\pm$02.26 91.03$\pm$02.03
5.5 In Depth Analysis
In this section, we delve into the reasons behind SAM’s poor performance on the COCO dataset, exploring two main hypotheses:
  • Decoder Limitations: The primary suspect for SAM’s subpar performance is the dependency of the image encoder on the prompt encoder and the mask decoder. To address this, we replaced the original mask decoder with a transformer decoder, as detailed in [15], modifying it for few-shot semantic segmentation following the inspiration from [41]. We introduced a constant prompt to the decoder, altering tokens to produce an $N$-class depth map rather than binary masks, and eliminated the IoU score output branch. We experimented with two mask decoder configurations, one with a Conv Transpose module and another where we upscaled the output through interpolation, while initializing remaining modules with pretrained weights. Despite these adjustments, this approach was outperformed by a simpler linear decoding method by up to 5% mIoU, likely due to the challenges of training a complex decoder in a Few-shot scenario.
  • Mask Distribution Bias: Another investigated issue was the mask distribution bias in COCO towards image centers, unlike the distribution in the SA-1B dataset used for SAM, highlighted in [15], where masks are more evenly distributed. We attempted to mitigate this bias with various preprocessing techniques (RandomCrop, CenterCrop, dividing each image into four smaller images); however, these modifications did not enhance our results.
These findings shed light on the specific challenges faced by SAM in adapting to the COCO dataset and highlight the complexities involved in optimizing for few-shot semantic segmentation tasks.
@@ 6.2 Semantic Segmentation and Few-Shot Semantic Segmentation
Few-shot semantic segmentation is not a recent topic. Earlier works, such as [31] and [26], explored multiple benchmarks, including PASCAL-$5^{i}$ and COCO-$20^{i}$, reminiscent of in-domain benchmarks used in few-shot classification. Subsequent research ventured into using pretrained models on ImageNet [39], resembling the Meta-dataset approach in few-shot classification. Notably, most of these studies assumed binary labelling for the few-shots, emphasizing a single object of interest in the available shot images, considering all other objects as background.
In contrast, our approach considers that the provided few-shots are fully labelled, and we leverage pretrained foundation models, not limited to those trained solely on ImageNet. This approach aligns more closely with contemporary industrial requirements, focusing on the utilization of the most efficient available models, without specific explicit restrictions.
@@ 7 Conclusion
In conclusion, DINO V2 emerges as the top-performing backbone model, consistently outperforming other foundation models. While SAM exhibits strong overall results, it presents unexpected limitations on certain datasets. Surprisingly, our experiments underscore the competitive advantages of straightforward methods such as simple linear segmentation heads trained on the few given shots, challenging the necessity of complex procedures.
This work not only presents a comparison of techniques but also a practical benchmark. We hope this benchmark will contribute to improving few-shot semantic segmentation methods and will provide a solid ground for comparison.